OpenStudy (anonymous):

A total of 60 gallons of gas is alloted to 2 vehicles. One vehicle is to receieve 12 gallons less than the other. How many gallons will each recieve? Please Help :)))

OpenStudy (anonymous):

1st vehicle = (a) 2nd vehicle = (a-12) (a-12) + a = 60 2a-12=60 2a=72 a= 36 so first vehicles is 36 2nd vehicle =36-12 so 2nd vehicle = 24

OpenStudy (anonymous):

Thanks